OpenStudy (anonymous):

2,7,4, (?), 8, 32, 16, 52

OpenStudy (anonymous):

the answer is .................. 17

OpenStudy (anonymous):

i think this is a complicated sequence 2+2=4 4+4=8 8+8=16 is the series of alternating terms and 7+10=17 17+15=32 32+20=52 is the way how other terms alternatively proceed

OpenStudy (anonymous):

yes thnaks